# Leveraging Perceptual Scores for Dataset Pruning in Computer Vision Tasks

### v1

@@ 2 Method
Another method for estimating the bits-per-pixel of an image is the log likelihood of that image inferred from a trained generative model [12]. In the past likelihood has been used for out-of-distribution detection, but as far as we know not for data pruning. We use it as a score, $NLL$, to prune data for the image classification task. It has also been observed [20] that generative likelihood of an image is inversely correlated with complexity of the image that is measured by the entropy of a lossless encoder. Subtracting the entropy from log likelihood of an image compensates for the complexity of the image [20]. This score, $CPX=NLL-BPP_{J}$ gives, surprisingly, the best results for the image classification task.
In  [21] authors cluster features, inferred from a self-supervised model, of images, using k-means algorithm. The distance to the nearest cluster centroid is the prototypicality score of an image, $PS$. k-means is a method for vector quantization and the distance to the nearest centroid is the distortion incurred in using the quantizer. By rate-distortion theory [2] we can consider the distortion to have a bit-per-pixel interpretation under lossy compression conditions. The difference between $PS$ and $BPP_{J}$ is that the latter uses JPEG which has no learning component. $PS$, on the other hand has a learning component but it is trained on a different dataset in an unsupervised manner. Also the distortion in $PS$ is on the decoded features, and in $BPP_{J}$ is on the decoded image; in fact its not clear that the features used in $PS$ can be used to decode (generate) the image. On the other hand $NLL$ and $CPX$ use generative model learnt on the dataset itself, and have the capability to generate (decode) the image itself. They are unsupervised generative models. Note that in this work we do not compare with the any other scores for coreset selection, data pruning, because they rely on supervised learning.
@@ 3 Results
Table 1: Accuracy (%) for CIFAR dataset for different sampling methods (rows) and different pruned sample size (columns). $RND$ uses no random sampling, $BPP_{J}$, $PS$, $NLL$, $CPX$ use the score only, while $BPP_{J}+G_{S}$ uses score and graph based sampling. Best results are obtained by using ascending order for all scores.
In this paper we use CIFAR-10 dataset for image classification experiments. We use Resent18 model with hyper-parameters as in [21]. For semantic segmentation we use VOC and ADE20K dataset, MobileNet model as the encoder, and one convolution along with deep supervision as the decoder  [26]. For ADE20K we use the default hyper-parameters, for VOC we lower the learning rate. We estimate the likelihood, $NLL$, of an image in CIFAR, using GLOW models [8, 12] trained on CIFAR. We use $NLL$ and $BPP_{J}$ to calculate $CPX$.
Table 1 shows the result for image classification task. $BPP_{J}$ does not do well at all, and neither does $NLL$. We tried sampling both in descending and ascending order of $BPP_{J}$ and $NLL$ and the best results we got were with descending order – that is with high bits-per-pixel images. Accounting for complexity of image, $CPX$, estimated by generative model does well, beating $PS$ and random sampling for lower pruning rates. Also using graph sampling substantially improves performance for $BPP_{J}$ score, showing the importance of data diversity.
In Fig. 4 and Fig. 5 semantic segmentation results for VOC and ADE dataset are shown. For both datasets we see that $BPP_{J}$ (again in descending order) by itself (without graph sampling) does better than prototypicality score (in ascending order). $BPP_{J}$ along with k-NN graph defined by histograms of labels $G_{H}$ give measurable improvement over random sampling and over graph sampling using SWAV featuers $G_{S}$. $PS$ does poorly for semantic segmentation because it exaggerates the class imbalance problem.
@@
In this paper we proposed a simple, intrinsic, perceptual complexity score for coreset selection. This score requires no training and does not use the labels; when we use labels, in $G_{H}$, we are not training on them and coreset selection can still be done before training starts. One of our original hypothesis of learning from iconic images does not seem, as of now, correct for deep learning models. We also believe that perceptual scores could be used a priors to be updated with label and task information, once training starts. The positive results of this study are usage of $CPX$ and on semantic segmenation.

### v2

@@ 2 Method
Another method for estimating the bits-per-pixel of an image is the log likelihood of that image inferred from a trained generative model [12]. In the past likelihood has been used for out-of-distribution detection, but as far as we know not for data pruning. We use it as a score, NLLNLL, to prune data for the image classification task. It has also been observed [20] that generative likelihood of an image is inversely correlated with complexity of the image that is measured by the entropy of a lossless encoder. Subtracting the entropy from log likelihood of an image compensates for the complexity of the image [20]. This score, $CPX=NLL-BPP_{J}$ gives, surprisingly, the best results for the image classification task.
In  [21] authors cluster features, inferred from a self-supervised model, of images, using k-means algorithm. The distance to the nearest cluster centroid is the prototypicality score of an image, PSPS. k-means is a method for vector quantization and the distance to the nearest centroid is the distortion incurred in using the quantizer. By rate-distortion theory [2] we can consider the distortion to have a bit-per-pixel interpretation under lossy compression conditions. The difference between PSPS and $BPP_{J}$ is that the latter uses JPEG which has no learning component. PSPS, on the other hand has a learning component but it is trained on a different dataset in an unsupervised manner. Also the distortion in PSPS is on the decoded features, and in $BPP_{J}$ is on the decoded image; in fact its not clear that the features used in PSPS can be used to decode (generate) the image. On the other hand NLLNLL and CPXCPX use generative model learnt on the dataset itself, and have the capability to generate (decode) the image itself. They are unsupervised generative models. Note that in this work we do not compare with the any other scores for coreset selection, data pruning, because they rely on supervised learning.
@@ 3 Results
Table 1: Accuracy (%) for CIFAR dataset for different sampling methods (rows) and different pruned sample size (columns). RNDRND uses no random sampling, $BPP_{J}$, PSPS, NLLNLL, CPXCPX use the score only, while $BPP_{J}+G_{S}$ uses score and graph based sampling. Best results are obtained by using ascending order for all scores.
In this paper we use CIFAR-10 dataset for image classification experiments. We use Resent18 model with hyper-parameters as in [21]. For semantic segmentation we use VOC and ADE20K dataset, MobileNet model as the encoder, and one convolution along with deep supervision as the decoder  [26]. For ADE20K we use the default hyper-parameters, for VOC we lower the learning rate. We estimate the likelihood, NLLNLL, of an image in CIFAR, using GLOW models [8, 12] trained on CIFAR. We use NLLNLL and $BPP_{J}$ to calculate CPXCPX.
Table 1 shows the result for image classification task. $BPP_{J}$ does not do well at all, and neither does NLLNLL. We tried sampling both in descending and ascending order of $BPP_{J}$ and NLLNLL and the best results we got were with descending order – that is with high bits-per-pixel images. Accounting for complexity of image, CPXCPX, estimated by generative model does well, beating PSPS and random sampling for lower pruning rates. Also using graph sampling substantially improves performance for $BPP_{J}$ score, showing the importance of data diversity.
In Fig. 4 and Fig. 5 semantic segmentation results for VOC and ADE dataset are shown. For both datasets we see that $BPP_{J}$ (again in descending order) by itself (without graph sampling) does better than prototypicality score (in ascending order). $BPP_{J}$ along with k-NN graph defined by histograms of labels $G_{H}$ give measurable improvement over random sampling and over graph sampling using SWAV featuers $G_{S}$. PSPS does poorly for semantic segmentation because it exaggerates the class imbalance problem.
@@
In this paper we proposed a simple, intrinsic, perceptual complexity score for coreset selection. This score requires no training and does not use the labels; when we use labels, in $G_{H}$, we are not training on them and coreset selection can still be done before training starts. One of our original hypothesis of learning from iconic images does not seem, as of now, correct for deep learning models. We also believe that perceptual scores could be used a priors to be updated with label and task information, once training starts. The positive results of this study are usage of CPXCPX and on semantic segmenation.